# Multi-objective Optimization in CPU Design Space Exploration: Attention is All You Need

### v1

@@ 7.1.2. Comparison on Best Balanced Designs
7.1.3. Comparison on Single-objective Predictor
An accurate predictor is fundamental to an effective DSE framework. Fig. 10 illustrates the comparison of the prediction models with the SOTA DSE frameworks. Fig. 10(a), (b), and (c) present the MAPE of AttentionDSE compared with three DSE frameworks. For example, AttentionDSE outperforms all workloads, achieving 37.1%, 33.4%, and 31.6% on average for IPC, Power, and Area prediction on MAPE, respectively, compared to the SOTA DSE framework (MoDSE). These results highlight the superiority of AttentionDSE over statistical regression methods. We also conduct experiments using an MLP regression model, as shown in Table 3, to demonstrate the superiority of the attention mechanism.
Table 3. Comparison of AttentionDSE with the MLP model.
@@ 7.1.3. Comparison on Single-objective Predictor
7.2. Optimization Effect Analysis
@@
7.2.1. Efficiency of Perception-driven Attention Mechanism
@@
Figure 11. The efficiency of PDA mechanism.
Fig. 11 shows the experiment results, the line chart illustrates the normalized training time, while the bar chart depicts the memory footprint. The training time for models without the PDA mechanism increases exponentially with the size of the design space. For example, when the design space expands sixfold, the training time increases by 30$\times$. In contrast, AttentionDSE with the PDA mechanism exhibits relatively stable memory usage as the number of parameters grows. This stability extends to large design spaces because the PDA mechanism only calculates within clusters that contain the most relevant parameters. Regarding memory footprint, in smaller design spaces, memory usage is primarily influenced by fixed data occupation during training. However, attention weights and other learnable parameters in larger design spaces gradually dominate memory usage.
Figure 12. The impact of the PDA mechanism.
Fig. 12 shows the difference in MAPE introduced by the PDA mechanism. Due to the PDA mechanism clustering the relevant parameters, captured by the sliding window attention technique, the predictive accuracy does not decrease and may even slightly increase. This improvement is primarily due to the enhanced understanding of local information within the pipeline stage provided by PDA. By focusing on parameter interactions within a defined window, PDA captures intricate relationships between parameters, resulting in more accurate predictions.
7.2.2. Analysis on Attention-aware Bottleneck Analysis Algorithm
The superiority of AttentionDSE derives from both its high-accuracy prediction model and the ABA algorithm. To more effectively illustrate the impact of the ABA algorithm and avoid interference from the inaccurate prediction model, we conduct experiments on AttentionDSE by substituting the ABA algorithm with other design point search methods. Due to the large design space in our experiment, exhaustive search methods like MoDSE become impractical. Additionally, methods similar to BOOMExplorer can only be applied to the GPR model. Therefore, we select the random sample search method as the baseline for comparison with the ABA algorithm. Fig. 13 visualizes the PHV curves and ADRS curves across iterations, comparing them with the random sample search method.
Figure 13. The convergence of ABA algorithm for IPC-Power optimization.
Regarding PHV convergence, the ABA algorithm notably achieves a higher PHV early on, consistently surpassing the random sample search method throughout various simulation iterations. Specifically, the ABA algorithm reaches a PHV of 38.2 within one hundred iterations, while the random sample search method requires more than 5$\times$ as many iterations to achieve similar levels. This efficiency results from the ABA algorithm’s effective analysis of parameter interactions and their impact on performance metrics, significantly reducing exploration time by 80% and guiding optimal parameter updates for enhanced performance.
@@ 7.2.2. Analysis on Attention-aware Bottleneck Analysis Algorithm
8. Conclusion
In this paper, we introduced AttentionDSE, a comprehensive DSE framework tailored for high-dimensional design spaces. AttentionDSE seamlessly synergizes the surrogate model with the acquisition function by incorporating the attention mechanism. The innovative PDA mechanism enhances prediction efficiency, while the ABA algorithm offers detailed insights into individual micro-architectural parameter contributions. These advancements improve the design process’s interpretability and minimize the need for expert knowledge. AttentionDSE represents the first attempt to apply attention mechanisms to DSE tasks, showcasing the potential of this approach in optimizing complex design spaces.

### v2

@@ VII-A2 Comparison on Best Balanced Designs
VII-A3 Comparison on Single-objective Predictor
An accurate predictor is fundamental to an effective DSE framework. Fig. 10 illustrates the comparison of the prediction models with the SOTA DSE frameworks. Fig. 10(a), (b), and (c) present the MAPE of AttentionDSE compared with three DSE frameworks. All the predictors are trained with one thousand design points. For example, AttentionDSE outperforms all workloads, achieving 37.1%, 33.4%, and 31.6% on average for IPC, Power, and Area prediction on MAPE, respectively, compared to the SOTA DSE framework (MoDSE). These results highlight the superiority of AttentionDSE over statistical regression methods. We also conduct experiments using an MLP regression model, as shown in Table III, to demonstrate the superiority of the attention mechanism.
TABLE III: Comparison of AttentionDSE with the MLP model.
Methods
@@
VII-B Optimization Effect Analysis
@@
VII-B1 Efficiency of Perception-driven Attention Mechanism
@@
Figure 11: The efficiency of PDA mechanism.
[…]
Figure 12: The impact of the PDA mechanism.
Fig. 12 shows the difference in MAPE introduced by the PDA mechanism. Due to the PDA mechanism clustering the relevant parameters, captured by the sliding window attention technique, the predictive accuracy does not decrease and may even slightly increase. This improvement is primarily due to the enhanced understanding of local information within the pipeline stage provided by PDA. By focusing on parameter interactions within a defined window, PDA captures relationships between parameters, improving prediction accuracy.
VII-B2 Analysis on Attention-aware Bottleneck Analysis Algorithm
The superiority of AttentionDSE derives from both its high-accuracy prediction model and the ABA algorithm. To more effectively illustrate the impact of the ABA algorithm and avoid interference from the inaccurate prediction model, we conduct experiments on AttentionDSE by substituting the ABA algorithm with other design point search methods. Given the large design space, exhaustive methods like MoDSE are impractical. Since BOOMExplorer-like methods apply only to GPR models, we use random sample search as the baseline to compare with ABA. Fig. 13 shows PHV and ADRS curves over iterations against this baseline.
Figure 13: The convergence of ABA algorithm for IPC-Power optimization.
[…]
@@ VII-B2 Analysis on Attention-aware Bottleneck Analysis Algorithm
VIII Conclusion
In this paper, we propose AttentionDSE, the first attention-driven framework that unifies performance prediction and architectural exploration for high-dimensional CPU design. By leveraging transformer-based attention, it achieves accurate PPA estimation and exposes parameter-wise bottlenecks to guide efficient design refinement. Through PDA and ABA, AttentionDSE delivers faster convergence and better Pareto quality than prior methods. This work demonstrates the potential of attention mechanisms as a scalable and interpretable solution for automated micro-architecture design.